A car starts moving along a line first with acceleration a=5ms-2 starting from rest then uniformly and finally decelerating at same rate a and comes to rest.The total time of motion is 25s.The average speed during time is 20 ms-1. How long does particle move uniformly?
1
Expert's answer
2019-06-24T15:44:57-0400

The average speed can be calculated as total distance over total time:


"s_\text{av}=\frac{D}{T}."

The total time is the sum of time when the car accelerated, moved uniformly with speed "s_u" within the distance "d_u", and decelerated:


"T=\frac{s_u}{a}+\frac{d_u}{s_u}+\frac{s_u}{a}=\frac{2s_u}{a}+\frac{d_u}{s_u}."

The total distance the car traveled is:


"D=s_\text{av}T=\frac{s_u^2}{2a}+d_u+\frac{s_u^2}{2a}=\frac{s_u^2}{a}+d_u."

So, using the first expression, we have two equations to solve them for the speed and distance labeled with index u:

"500=\frac{s_u^2}{5}+d_u,"

"25=\frac{2s_u}{5}+\frac{d_u}{s_u}."

This will give you a quadratic equation for speed and the following pair of roots is the solution:


"s_u=25\text{ m/s},\space d_u=375\text{ m},\\
s_u=100\text{ m/s},\space d_u=-1500\text{ m}."

Both roots give the same time of uniform motion:


"t_u=\frac{d_u}{s_u}=15\text{ s}."